OpenStudy (anonymous):

A bat and a ball cost one dollar and ten cents in total. The bat costs a dollar more than the ball. How much does the ball cost? /does anyone know?

OpenStudy (kate_kat):

umm

OpenStudy (kate_kat):

10 cents?

OpenStudy (anonymous):

thats what i thought but it was wrong

OpenStudy (kate_kat):

hmm...

OpenStudy (tennis5518):

......?

OpenStudy (anonymous):

srsly no one knows?

OpenStudy (tennis5518):

I do what do you think?

OpenStudy (anonymous):

i said 10cents

OpenStudy (raffle_snaffle):

well 10 cents is wrong XD

OpenStudy (kate_kat):

i think it can range from 1cents to 9cents

OpenStudy (anonymous):

thx kate

OpenStudy (jdoe0001):

well... can you post a screenshot of the material, so we can see what's in it? as it stand, it IS 10 cents

OpenStudy (kate_kat):

:)

OpenStudy (raffle_snaffle):

Oooooo

OpenStudy (anonymous):

???

OpenStudy (anonymous):

lol its not 10cents

OpenStudy (jdoe0001):

hmmm

OpenStudy (kate_kat):

lol

OpenStudy (jdoe0001):

hmmm ok, I see the issue, hold the mayo

OpenStudy (anonymous):

kate said already, its like 5cents

OpenStudy (tennis5518):

i think 5cent

OpenStudy (kate_kat):

lol, yay

OpenStudy (jdoe0001):

so... say... the ball is "a", and the bat is "b" we know that a+b = 110 cents or $1.10 so a+ b = 110 but we also know that, the bat is 100 cents more than the ball, or "a" thus b = a + 100 so... \(\bf a+b=110\qquad \qquad {\color{brown}{ b=a+100}} \\ \quad \\ a+{\color{brown}{ a+100}}=110\implies 2a+100=110\) pretty sure you can take it from there

OpenStudy (kate_kat):

yeah thats what i said

OpenStudy (anonymous):

ok im gonna close the question

OpenStudy (tennis5518):

But the real answer is 5¢.

OpenStudy (anonymous):

ok...

OpenStudy (tennis5518):

yes...?

OpenStudy (anonymous):

bye everyone